# Modeling Cell Dynamics and Interactions with Unbalanced Mean Field Schrödinger Bridge

### v2

@@ 5.2 Reformulating the Loss with Weighted Particle Simulation
6 Results
Next, we evaluate CytoBridge’s ability to simultaneously learn cell dynamics and cell-cell interactions. In computations below, we take $\Psi(g(\mathbf{x},t))=g^{2}(\mathbf{x},t)$ and $\sigma(t)$ is constant. We also assume the interaction term is dependent on the distance between cells in gene expression space and we use radial basis functions (RBFs) to approximate it (Section A.3).
Synthetic Gene Regulatory Network
In order to examine CytoBridge’s capabilities of learning cell dynamics as well as their underlying interactions simultaneously, we conducted experiments on the three-gene simulation model following (Zhang et al., 2025a). The dynamics of the original three-gene model are governed by stochastic ordinary differential equations, incorporating self-activation, mutual inhibition, and external activation (Section B.1), as shown in Fig. 2 (a). Additionally, we incorporated interactions into the simulation process. We consider the following types of interactions: (1) attractive interactions. (2) Lennard-Jones-like potential (both attractive and repulsive) (3) no interactions. We aim to test whether CytoBridge can recover interaction in each case. For attractive interactions, the cells with similar gene expressions tend to converge toward similar levels. The interaction potential $\Phi$ is defined as: $\Phi(\mathbf{x}-\mathbf{y})=\left\|\mathbf{x}-\mathbf{y}\right\|^{2}$ As shown in Fig. 2(b), the dynamics of the three-gene model exhibit a quiescent area as well as an area with notable transition and increasing cell numbers. Moreover, the attractive potential results in the reduction in variance of the observed data at different time points. We compared CytoBridge with other methods across all time points and summarized the results ($\mathcal{W}_{1}$ and TMV) in Table 1. It is shown that CytoBridge achieves the best performance in both distribution matching and mass matching. Balanced Schrödinger bridge (e.g., (Tong et al., 2024b)), which neglects both growth and interaction terms, results in false transition and variance patterns (Fig. 2(c)). DeepRUOT (Zhang et al., 2025a), which is an unbalanced Schrödinger Bridge solver, exhibits correct transition patterns but fails to capture the reduction in variance as it neglects the cell-cell interactions (Fig. 2(d)). By leveraging the UMFSB framework, as shown in Fig. 2(e), CytoBridge correctly models both the transition patterns and the reduction in variance as the correct interaction potential (Fig. 2(g)) and growth rate (Fig. 2(f)) can be directly learned by CytoBridge. Also, CytoBridge is capable of constructing the underlying Waddington landscape (Fig. 2(h)). Other than the attractive interaction potential, we also incorporated the Lennard-Jones-like potential, and no interaction case. Results can be found in Section B.1, Figs. 4 and 5 and we find CytoBridge can correctly identify both the LJ potential and no interaction in these cases. Overall, both quantitative results and qualitative analysis indicate the necessity of incorporating both growth and interaction terms.
Figure 2: (a) Illustration of the synthetic gene regulatory dynamics. (b) The ground truth cellular dynamics project on $(X_{1},X_{2})$. (c) The dynamics learned by balanced Schrödinger bridge SF2M (Tong et al., 2024b). (d) The dynamics learned by DeepRUOT. (e) The dynamics learned by CytoBridge. (f) The learned interaction potential. (g) The growth rates inferred by CytoBridge. (h) The constructed landscape at $t=4$.
$t=1$ $t=2$ $t=3$ $t=4$
SF2M (Tong et al., 2024b) 0.146$\pm$0.002 0.080$\pm$0.000 0.320$\pm$0.004 0.250$\pm$0.000 0.447$\pm$0.005 0.515$\pm$0.000 0.554$\pm$0.005 0.930$\pm$0.000
Meta FM (Atanackovic et al., 2025) 0.374$\pm$0.000 0.080$\pm$0.000 0.414$\pm$0.000 0.250$\pm$0.000 0.496$\pm$0.000 0.515$\pm$0.000 0.479$\pm$0.000 0.930$\pm$0.000
MMFM (Rohbeck et al., 2025) 0.686$\pm$0.000 0.080$\pm$0.000 0.599$\pm$0.000 0.250$\pm$0.000 0.499$\pm$0.000 0.515$\pm$0.000 0.407$\pm$0.000 0.930$\pm$0.000
MIOFlow (Huguet et al., 2022) 0.315$\pm$0.000 0.080$\pm$0.000 0.387$\pm$0.000 0.250$\pm$0.000 0.483$\pm$0.000 0.515$\pm$0.000 0.518$\pm$0.000 0.930$\pm$0.000
uAM (Neklyudov et al., 2023) 0.489$\pm$0.000 0.081$\pm$0.000 0.995$\pm$0.000 0.033$\pm$0.000 1.402$\pm$0.000 0.459$\pm$0.000 1.655$\pm$0.000 1.516$\pm$0.000
UDSB (Pariset et al., 2023) 1.131$\pm$0.009 0.018$\pm$0.006 1.489$\pm$0.018 0.135$\pm$0.014 1.455$\pm$0.022 0.447$\pm$0.011 0.543$\pm$0.015 1.018$\pm$0.035
TIGON (Sha et al., 2024) 0.169$\pm$0.000 0.097$\pm$0.000 0.184$\pm$0.000 0.165$\pm$0.000 0.167$\pm$0.000 0.210$\pm$0.000 0.179$\pm$0.000 0.384$\pm$0.000
DeepRUOT (Zhang et al., 2025a) 0.044$\pm$0.002 0.014$\pm$0.007 0.045$\pm$0.002 0.026$\pm$0.018 0.053$\pm$0.002 0.059$\pm$0.032 0.057$\pm$0.003 0.075$\pm$0.044
CytoBridge (Ours) 0.015$\pm$0.001 0.013$\pm$0.009 0.014$\pm$0.001 0.021$\pm$0.024 0.018$\pm$0.002 0.043$\pm$0.041 0.038$\pm$0.003 0.058$\pm$0.061
@@ 6 Results
To demonstrate the scalability of CytoBridge to high-dimensional data, we adopt the mouse hematopoiesis dataset (Weinreb et al., 2020) which includes 49,302 cells with lineage tracing data collected at three time points. We use PCA to reduce the dimensions to 50 and serve as the input of CytoBridge. The dataset comprises diverse cell states and demonstrates pronounced cell division. Consequently, accurately modeling both cellular dynamics and growth rates is critical for reliable inference of cell fate. As shown in Table 2, CytoBridge outperforms other state-of-the-art methods in distribution matching, highlighting CytoBridge’s capabilities of capturing transition patterns (Fig. 3(a)). Besides, evidenced by the TMV metric, CytoBridge is able to recover the increase in cell numbers by learning the growth rate (Fig. 3(b)). The score function learned by CytoBridge indicates the existence of multiple attractors, which may lead to different cell fates (Fig. 3(c)). As shown in Fig. 3(d), The interactions learned by CytoBridge indicate that cells may tend to attract others with similar gene expression profiles. To further demonstrate the impact of cell-cell interactions on the transition of cells, we computed the correlation of each cell’s drift and interacting force. As shown in Fig. 9(a), the learned cell-cell interactions may promote early-stage cell differentiation and inhibit later-stage cell differentiation. Some additional results can be found in Section B.2.
Figure 3: Application in mouse blood hematopoiesis data ($\sigma=0.1$). (a) The overall velocity learned by CytoBridge. (b) The growth rates learned by CytoBridge. (c) The score function learned by CytoBridge at $t=2$. (d) The interactions inferred by CytoBridge.
Table 2: Wasserstein distance ($\mathcal{W}_{1}$) and Total Mass Variation (TMV) of predictions at different time points across five runs on mouse hematopoiesis data ($\sigma=0.1$). We show the mean value with one standard deviation, where bold indicates our algorithm as the best and $\underline{\text{underline}}$ indicates the second-best among all algorithms.
@@
SF2M (Tong et al., 2024b) 8.324$\pm$0.004 2.231$\pm$0.000 14.190$\pm$0.010 5.399$\pm$0.000
Meta FM (Atanackovic et al., 2025) 9.303$\pm$0.000 2.231$\pm$0.000 13.316$\pm$0.000 5.399$\pm$0.000
MMFM (Rohbeck et al., 2025) 9.098$\pm$0.000 2.231$\pm$0.000 12.167$\pm$0.000 5.399$\pm$0.000
MIOFlow (Huguet et al., 2022) 6.313$\pm$0.000 2.231$\pm$0.000 6.746$\pm$0.000 5.399$\pm$0.000
uAM (Neklyudov et al., 2023) 7.537$\pm$0.000 2.875$\pm$0.000 9.762$\pm$0.000 5.670$\pm$0.000
UDSB (Pariset et al., 2023) 10.687$\pm$0.058 0.282$\pm$0.146 13.477$\pm$0.053 3.010$\pm$0.225
TIGON (Sha et al., 2024) 6.140$\pm$0.000 1.234$\pm$0.000 6.973$\pm$0.000 2.083$\pm$0.000
DeepRUOT (Zhang et al., 2025a) 6.052$\pm$0.002 0.200$\pm$0.001 6.757$\pm$0.006 0.260$\pm$0.007
CytoBridge (Ours) 6.013$\pm$0.002 0.208$\pm$0.001 6.644$\pm$0.011 0.078$\pm$0.013
@@
We further evaluated CytoBridge on the Embryoid Body single-cell data which consists of 16,819 cells collected at five time points (Moon et al., 2019). We use PCA to reduce the dimensions to 50 and serve as the input of CytoBridge. As shown in Table 8, CytoBridge generally outperforms other methods in distribution matching and maintains mass-matching results comparable to other unbalanced algorithms. We plotted the learned velocity and score in Fig. 6, indicating the different cell fates learned by CytoBridge. The correlation of each cell’s drift and interacting force is shown in Fig. 9(b) indicates that the learned cell-cell interactions may promote cell differentiation while inhibiting cell differentiation of some outliers. We further consider the 3D-cultured in vitro pancreatic $\beta$-cell differentiation dataset (Veres et al., 2019) and A549 lung cancer cell line epithelial-mesenchymal transition (EMT) dataset induced by TGFB1 (Cook & Vanderhyden, 2020). We interestingly find that the interaction in the A549 cell line EMT process may be very weak. The detailed results can be found in Sections B.3, B.4 and B.5.
@@ 7 Conclusion
We have introduced CytoBridge for learning unbalanced stochastic mean-field dynamics from time-series snapshot data. CytoBridge transforms the SDE constraints into the ODE leveraging Fisher regularization for more efficient simulation in training. We have demonstrated the effectiveness of our method on both synthetic gene regulatory networks and single-cell RNA-seq data, showing its promising performance. We discuss limitations of current work and aspects for further exploration in Appendix E. Overall, CytoBridge provides a unified framework for generative modeling of time-series transcriptomics data, enabling more robust and realistic inference of underlying biological dynamics.
Acknowledgements
We thank Dr. Zhenfu Wang (PKU), Prof. Chao Tang (PKU) and Prof. Qing Nie (UCI) for their insightful discussions. This work was supported by the National Key R&D Program of China (No. 2021YFA1003301 to T.L.) and National Natural Science Foundation of China (NSFC No. 12288101 to T.L. & P.Z., and 8206100646, T2321001 to P.Z.). We acknowledge the support from the High-performance Computing Platform of Peking University for computation.
@@ A.3 Modeling Interaction Potential
The RBF centers $\mu_{k}$ are initialized by uniformly discretizing the exponentially transformed distance interval that corresponds to original distances from 00 to a predefined cutoff $d_{\text{cutoff}}$. The width parameters $\beta_{k}$ are initialized as the inverse square of a term proportional to the average spread of these centers in the transformed exponential scale.
@@
$k(\mathbf{x}_{i},\mathbf{x}_{j})=\begin{cases}1&\text{if }d_{ij}<d_{\text{cutoff}}\\ 0&\text{if }d_{ij}\geq d_{\text{cutoff}}\end{cases}$
@@ A.4 Reconstruction Loss Function
Here, $w_{i}(t_{k})$ is the weight of the $i$-th sampled particle at time $t_{k}$, $n_{k}$ denotes the number of cells in the original dataset at time $t_{k}$. The term $\text{card}\left(h_{k}^{-1}(\mathbf{x}_{i}(t_{k}))\right)$ represents the number of sampled data points from $A_{k}$ that are mapped to the $i$-th sampled particle $\mathbf{x}_{i}(t_{k})$ via the mapping $h_{k}$. This mapping $h_{k}$ assigns each real data point in $A_{k}$ to its closest sampled particle in $\hat{A}_{k}$. Essentially, the formula measures the squared difference between the particle’s current weight and the proportion of real data points it represents. Thus, it provides a fine-grained guidance on the weights of particles. Moreover, under the condition that $M_{k}=0$, it follows that $\sum_{i=1}^{N}w_{i}(t_{k})=\frac{1}{N}\sum_{i=1}^{N}\text{card}\left(h_{k}^{-1}(\mathbf{x}_{i}(t_{k}))\right)\frac{n_{k}}{n_{0}}=n_{k}/n_{0}$. Therefore the local matching loss also encourages the alignment of the total mass. The local mass matching loss $\mathcal{L}_{\text{Local Mass}}$ is then calculated by summing these errors over all time steps from $k=1$ to $T-1$: $\mathcal{L}_{\text{Local Mass}}=\sum_{k=1}^{T-1}M_{k}$.
@@
Here, The global mass matching loss $\mathcal{L}_{\text{Global Mass}}$ is then calculated by summing these errors over all time steps from $k=1$ to $T-1$: $\mathcal{L}_{\text{Global Mass}}=\sum_{k=1}^{T-1}G_{k}$.
@@ B.1 Synthetic Gene Regulatory Network
$dt$ 1 Time step size
@@
To evaluate CytoBridge’s capability of recovering distributions at unseen time points, we conducted hold-one-out experiments on synthetic gene regulatory data with attractive interactions. Specifically, intermediate time points were individually left out during training and subsequently recovered during evaluation. Note that the default setting of UDSB only supports data with three time points as inputs, we only compare with the performance of other methods. As shown in Table 4, CytoBridge performed best on all held-out time points, indicating its ability to infer reasonable trajectories from snapshots.
@@
SF2M (Tong et al., 2024b) 0.638$\pm$0.005 0.466$\pm$0.007 0.507$\pm$0.007
Meta FM (Atanackovic et al., 2025) 1.232$\pm$0.000 1.371$\pm$0.000 2.732$\pm$0.000
MMFM (Rohbeck et al., 2025) 0.476$\pm$0.000 0.776$\pm$0.000 1.825$\pm$0.000
MIOFlow (Huguet et al., 2022) 0.225$\pm$0.000 0.270$\pm$0.000 0.234$\pm$0.000
uAM (Neklyudov et al., 2023) 0.600$\pm$0.000 0.975$\pm$0.000 1.243$\pm$0.000
TIGON (Sha et al., 2024) 0.254$\pm$0.000 0.214$\pm$0.000 0.178$\pm$0.000
DeepRUOT (Zhang et al., 2025a) 0.184$\pm$0.001 0.086$\pm$0.004 0.079$\pm$0.003
CytoBridge (Ours) 0.182$\pm$0.001 0.064$\pm$0.004 0.043$\pm$0.002
@@
where $d_{e}$ represents the distance at which the repulsive and attractive forces balance each other. Moreover, to avoid the singularity of the potential function, we clipped the magnitude of forces to a pre-defined value $F_{max}$. As shown in Fig. 4, CytoBridge is able to reproduce the Lennard-Jones interaction potential with only information from snapshots provided, leading to correct transition and change of variance. We also conducted quantitative evaluations compared with DeepRUOT, in which the cellular interactions are neglected. As shown in Table 5, CytoBridge consistently outperforms DeepRUOT across all time points, underscoring the importance of explicitly modeling cellular interactions.
@@
Despite the good performance of CytoBridge on the synthetic gene model with different types of interactions, it yet remains unknown whether our method can help identify the cases where the ground truth dynamics itself does not involve cellular interactions. Therefore, we further conduct experiments on the synthetic gene model without interactions involved by setting $\Phi(\mathbf{x}-\mathbf{y})=0$. As shown in Table 6, explicitly incorporating cellular interactions does not result in the improvement of performance in distribution matching at most time points. The main reason for this phenomenon is that it may be hard for a neural network to directly infer all-zero outputs. The numerical error resulting from our interaction network may lead to perturbations at certain time points and thereby exhibit poorer performance than methods that neglect the interaction term. We further analyze the learned interaction forces acting on each cell. As shown in Fig. 5, compared with forces learned from the dynamics with explicit interactions, forces learned from dynamics with no interaction exhibit notably smaller magnitudes. We compute the correlation of learned velocity and forces of each cell to examine whether the learned forces from data without interaction may exhibit certain patterns related to the transitions of cells. Moran’s I, which is a statistic that measures spatial autocorrelation, indicating the degree to which nearby locations have similar values, is utilized to identify whether these patterns exist. Moran’s I calculated from data with ground truth interactions is 0.514, while that calculated from data without interactions is 0.281, which is notably smaller. This may indicate that instead of explainable correlated patterns between transition and interaction, interacting forces learned from data without interaction exhibit random patterns to some extent. Biologically, it remains challenging to precisely know whether and how cells interact with each other. Based on this difficulty, we hope that our preceding observations can help explain the performance of our method when applied to real-world biological datasets. Furthermore, we aim for this analysis to provide valuable biological insight into the presence and nature of cell-cell interactions.

### v3

@@ 6 Results
[…]
@@
To demonstrate the scalability of CytoBridge to high-dimensional data, we adopt the mouse hematopoiesis dataset (Weinreb et al., 2020) which includes 49,302 cells with lineage tracing data collected at three time points. We use PCA to reduce the dimensions to 50 and serve as the input of CytoBridge. The dataset comprises diverse cell states and demonstrates pronounced cell division. Consequently, accurately modeling both cellular dynamics and growth rates is critical for reliable inference of cell fate. As shown in Table˜2, CytoBridge outperforms other state-of-the-art methods in distribution matching, highlighting CytoBridge’s capabilities of capturing transition patterns (Fig.˜3(a)). Besides, evidenced by the TMV metric, CytoBridge is able to recover the increase in cell numbers by learning the growth rate (Fig.˜3(b)). The regions with high learned growth rates correspond to the hematopoietic stem cell populations. This is biologically consistent with the lineage tracing barcode results (Sha et al., 2024). The score function learned by CytoBridge indicates the existence of multiple attractors, which may lead to different cell fates (Fig.˜3(c)). To further demonstrate the impact of cell-cell interactions on the transition of cells, we computed the correlation of each cell’s drift and interacting force. As shown in Fig.˜3(d), the learned cell-cell interactions may promote early-stage cell differentiation and inhibit later-stage cell differentiation. Some additional results can be found in Section˜B.2.
Figure 3: Application in mouse blood hematopoiesis data ($\sigma=0.1$), visualized in UMAP space. (a) The overall velocity learned by CytoBridge. (b) The growth rates learned by CytoBridge. (c) The score function learned by CytoBridge at $t=2$. (d) The correlation of velocity and interacting forces.
Table 2: Wasserstein distance ($\mathcal{W}_{1}$) and Total Mass Variation (TMV) of predictions at different time points across five runs on mouse hematopoiesis data ($\sigma=0.1$). We show the mean value with one standard deviation, where bold indicates our algorithm as the best among all algorithms.
@@
[…]
@@
We further evaluated CytoBridge on the Embryoid Body single-cell data which consists of 16,819 cells collected at five time points (Moon et al., 2019). We used PCA to reduce the dimensions to 50 and serve as the input of CytoBridge. As shown in Table˜9, CytoBridge generally outperforms other methods in distribution matching and maintains mass-matching results comparable to other unbalanced algorithms. We plotted the learned velocity and score in Fig.˜6, indicating the different cell fates learned by CytoBridge. The correlation of each cell’s drift and interacting force is shown in Fig.˜6(d) indicates that the learned cell-cell interactions may promote cell differentiation while inhibiting cell differentiation of some outliers. We also consider the 3D-cultured in vitro pancreatic $\beta$-cell differentiation dataset (Veres et al., 2019) and A549 lung cancer cell line epithelial-mesenchymal transition (EMT) dataset induced by TGFB1 (Cook and Vanderhyden, 2020). Interestingly, we find that the interaction in the A549 cell line EMT process may be very weak. The detailed results can be found in Sections˜B.3, B.4 and B.5.
Extension to Spatiotemporal Transcriptomics
To demonstrate CytoBridge’s applicability in modeling cellular interactions with explicit physical proximity, we applied CytoBridge to a zebrafish spatiotemporal transcriptomics dataset (Liu et al., 2022a), using slices from 5.25 hpf and 10 hpf as input. We evaluated the performance of CytoBridge on the task of reconstructing the dynamics of cell spatial migration, as well as gene expression, with a separate velocity and interactions for physical space and gene expression space respectively. As shown in Table˜13, CytoBridge outperforms other methods in both tasks. The results are visualized in Fig.˜9. Furthermore, downstream interpretability analysis of the learned interactions identified biologically relevant pathways crucial for zebrafish development. Detailed results can be found in Section˜B.6. The preliminary application of CytoBridge to spatiotemporal transcriptomics demonstrates our framework’s potential in modeling spatially resolved data.
@@ 7 Conclusion
We have introduced CytoBridge for learning unbalanced stochastic mean-field dynamics from time-series snapshot data. To tackle the interacting particle system inference from temporal snapshots, CytoBridge transforms the SDE constraints into ODE leveraging Fisher regularization for more efficient simulation in training. We have demonstrated the effectiveness of our method on both synthetic gene regulatory networks and single-cell RNA-seq data, showing its promising performance. Overall, CytoBridge provides a unified framework for generative modeling of time-series transcriptomics data, enabling more robust and realistic inference of underlying biological dynamics.
Limitations and Further Directions While CytoBridge offers valuable insights into incorporating cell-cell interaction with unbalanced stochastic dynamics, several aspects could benefit from further exploration. Firstly, CytoBridge minimizes the upper bound of the energy term in the UMFSB problem. Directly optimizing the original UMFSB formulation remains an important question. Potential solutions may involve leveraging neural SDEs or the integration-by-parts strategy proposed by (Zhang et al., 2025a) as viable approaches. Secondly, the current neural network parameterization of cellular interaction terms is based on techniques used to reduce degrees of freedom, such as RBF expansion. Future work could explore sparse representation methods to replace it for improved expressive power. Thirdly, the training of CytoBridge involves multiple stages. Simplifying the training process by incorporating optimality conditions (e.g., HJB equations) presents a promising research direction (Sun et al., 2025). Furthermore, the current modeling of cellular interactions have not incorporated certain biological priors such as ligand-receptor information. We plan to explore this direction in future work to further refine our approach. Finally, extending the concept of flow matching to this context and developing simulation-free training methods for stochastic, growth/death, and interaction dynamics could also advance the CytoBridge’s applicability.
Acknowledgments and Disclosure of Funding
We thank Dr. Zhenfu Wang (PKU), Prof. Chao Tang (PKU) and Prof. Qing Nie (UCI) for their insightful discussions. This work was supported by the National Key R&D Program of China (No. 2021YFA1003301 to T.L.), National Natural Science Foundation of China (NSFC No. 12288101 to T.L. & P.Z., and 8206100646, T2321001 to P.Z.) and The Fundamental Research Funds for the Central Universities, Peking University. We acknowledge the support from the High-performance Computing Platform of Peking University for computation.
@@ A.3 Modeling Interaction Potential
The RBF centers $\mu_{k}$ are initialized by uniformly discretizing the exponentially transformed distance interval that corresponds to original distances from 0 to a predefined cutoff $d_{\text{cutoff}}$. The width parameters $\beta_{k}$ are initialized as the inverse square of a term proportional to the average spread of these centers in the transformed exponential scale. The RBF expansion allows our model to learn interactions between different cells by encoding interaction distances across multiple scales, rather than enforcing interactions only between similar cells.
@@
[…]
@@ A.4 Reconstruction Loss Function
Here, $w_{i}(t_{k})$ is the weight of the $i$-th sampled particle at time $t_{k}$, $n_{k}$ denotes the number of cells in the original dataset at time $t_{k}$. The term $\text{card}\left(h_{k}^{-1}(\mathbf{x}_{i}(t_{k}))\right)$ represents the number of sampled data points from $A_{k}$ that are mapped to the $i$-th sampled particle $\mathbf{x}_{i}(t_{k})$ via the mapping $h_{k}$. This mapping $h_{k}$ assigns each real data point in $A_{k}$ to its closest sampled particle in $\hat{A}_{k}$. Essentially, the formula measures the squared difference between the particle’s current weight and the proportion of real data points it represents. Thus, it provides a fine-grained guidance on the weights of particles. Moreover, under the condition that $M_{k}=0$, it follows that $\sum_{i=1}^{N}w_{i}(t_{k})=\frac{1}{N}\sum_{i=1}^{N}\text{card}\left(h_{k}^{-1}(\mathbf{x}_{i}(t_{k}))\right)\frac{n_{k}}{n_{0}}=n_{k}/n_{0}$. Therefore the local matching loss also encourages the alignment of the total mass. The local mass matching loss $\mathcal{L}_{\text{Local Mass}}$ is then calculated by summing these errors over all time steps from $k=1$ to T1T-1: $\mathcal{L}_{\text{Local Mass}}=\sum_{k=1}^{T-1}M_{k}$.
@@
Here, The global mass matching loss $\mathcal{L}_{\text{Global Mass}}$ is then calculated by summing these errors over all time steps from $k=1$ to T1T-1: $\mathcal{L}_{\text{Global Mass}}=\sum_{k=1}^{T-1}G_{k}$.
@@ B.1 Synthetic Gene Regulatory Network
dtdt 1 Time step size
@@
To evaluate CytoBridge’s capability of recovering distributions at unseen time points, we conducted hold-one-out experiments on synthetic gene regulatory data with attractive interactions. Specifically, intermediate time points were individually left out during training and subsequently recovered during evaluation. Note that the default setting of UDSB only supports data with three time points as inputs, we only compare with the performance of other methods. As shown in Table˜4, CytoBridge performed best on all held-out time points, indicating its ability to infer reasonable trajectories from snapshots.
@@
[…]
@@
where $d_{e}$ represents the distance at which the repulsive and attractive forces balance each other. Moreover, to avoid the singularity of the potential function, we clipped the magnitude of forces to a pre-defined value $F_{max}$. As shown in Fig.˜4, CytoBridge is able to reproduce the Lennard-Jones interaction potential with only information from snapshots provided, leading to correct transition and change of variance. We also conducted quantitative evaluations compared with DeepRUOT, in which the cellular interactions are neglected. As shown in Table˜5, CytoBridge consistently outperforms DeepRUOT across all time points, underscoring the importance of explicitly modeling cellular interactions.
@@
Despite the good performance of CytoBridge on the synthetic gene model with different types of interactions, it yet remains unknown whether our method can help identify the cases where the ground truth dynamics itself does not involve cellular interactions. Therefore, we further conduct experiments on the synthetic gene model without interactions involved by setting $\Phi(\mathbf{x}-\mathbf{y})=0$. As shown in Table˜6, explicitly incorporating cellular interactions does not result in the improvement of performance in distribution matching at most time points. The main reason for this phenomenon is that it may be hard for a neural network to directly infer all-zero outputs. The numerical error resulting from our interaction network may lead to perturbations at certain time points and thereby exhibit poorer performance than methods that neglect the interaction term. We further analyze the learned interaction forces acting on each cell. As shown in Fig.˜5, compared with forces learned from the dynamics with explicit interactions, forces learned from dynamics with no interaction exhibit notably smaller magnitudes. We compute the correlation of learned velocity and forces of each cell to examine whether the learned forces from data without interaction may exhibit certain patterns related to the transitions of cells. Moran’s I, which is a statistic that measures spatial autocorrelation, indicating the degree to which nearby locations have similar values, is utilized to identify whether these patterns exist. Moran’s I calculated from data with ground truth interactions is 0.514, while that calculated from data without interactions is 0.281, which is notably smaller. This may indicate that instead of explainable correlated patterns between transition and interaction, interacting forces learned from data without interaction exhibit random patterns to some extent. Biologically, it remains challenging to precisely know whether and how cells interact with each other. Based on this difficulty, we hope that our preceding observations can help explain the performance of our method when applied to real-world biological datasets. Furthermore, we aim for this analysis to provide valuable biological insight into the presence and nature of cell-cell interactions.